40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 1 11:31:50 2011, josephb, Update, CDS, MC1 LR digital reading close to zero, readback ~0.7 volts 
    Reply  Sun Jun 26 22:38:20 2011, rana, Update, CDS, MC1 LR dead for > 1 month; now revived temporarily lrsen.pngbt.jpg
       Reply  Mon Jun 27 00:23:11 2011, rana, Update, CDS, ETMX SIDE problem etmx-side.png
       Reply  Fri Nov 18 01:12:14 2011, Suresh, Update, CDS, MC1 LR dead for > 1 month; now revived temporarily 
Message ID: 5938     Entry time: Fri Nov 18 01:12:14 2011     In reply to: 4888
Author: Suresh 
Type: Update 
Category: CDS 
Subject: MC1 LR dead for > 1 month; now revived temporarily 

[Den, Mirko, Suresh]

    We were investigating why there is no correlation between MC1 osem signals and seismic motion.   During this we noticed a recurrence of this old problem of MC1_LR sensor being dead.  I went and pressed down the chip holders where the AA filters used to sit and which now hold the jumper wire.  The board is large and flexible it is quite likely some solder joint is broken on the MC1_LR path on this board.

   The signal came back to life and is okay now. But it can break off again any time.




 Since the MC1 LRSEN channel is not wasn't working, my input matrix diagonalization hasn't worked today wasn't working. So I decided to fix it somehow.

I went to the rack and traced the signal: first at the LEMO monitor on the whitening card, secondly at the 4-pin LEMO cable which goes into the AA chassis.

The signal existed at the input to the AA chassis but not in the screen. So I pressed the jumper wire (used to be AA filter) down for the channel corresponding to the MC1 LRSEN channel.

It now has come back and looks like the other sensors. As you can see from this plot and Joe's entry from a couple weeks ago, this channel has been dead since May 17th.

The ELOG reveals that Kiwamu caught Steve doing some (un-elogged) fooling around there. Burnt Toast -> Steve.


993190663   =      free swinging ringdown restarted again


ELOG V3.1.3-